a wine merchant wishes to get 300 liters of wine that sells $3 per liter.She wishes to mix wines selling at $4.5,$1.5 and $2.25 per liter respectively.She must use twice as much of the $2.25 wine as the $1.5 wine.How many liters of each should she use?
Answer by mananth(16949)

$1.5 ----x
$2.25-----2x
$4.5------300-3x
...
1.5*x+2.25*2x+4.5*(300-3x)=300*3
6x-13.50x+1350=900
-7.5x=-450
/-7.5
x=60 liters------- $1.5
120liters----------$2.25
120liters----------$4.50